# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v2

@@ 4 Faster mLSTM with Sigmoid Input Gate
Figure 4: Transfer behavior of the mLSTM before and after the RMS-norm layer (ϵ=italic-$\epsilon=$1e-6) for different input and forget gate values. The color shows the gain of the mLSTM defined in (26). After the norm layer mLSTMexp and mLSTMsig exhibit the same transfer behavior.
The mLSTM with exponential gating (i.e. exponential input gate) introduced by Beck et al. (2024) requires to compute and keep track of two additional states, the normalizer state $\bm{n}_{t}$ and max state $m_{t}$, as we show in Appendix D.1.
Both will increase kernel runtime: The normalizer must be computed through summations, and tracking the max state throughout the tiled computation in TFLA (see Section 3.2 and C.1) prevents efficient fusing of loops within the kernel (see Appendix C.2).
Additionally, our analysis in Section 4.2 suggests to initialize the input gate biases at larger negative values (e.g. -10), such that the input gate pre-activations can grow slowly during training. We observe that most of these values stay below 0 during training (see Figure 14 in Appendix E). Therefore, we seek an alternative activation function which is similar to the exponential function in the negative range, but bounded in the positive range. This suggests to use the sigmoid function
$\sigma(x)=\frac{1}{1+\exp(-x)}=\frac{\exp(x)}{\exp(x)+1},$ (21)
which converges to $\exp(x)$ for $x\rightarrow-\infty$ and $1$ for $x\rightarrow\infty$.
@@
The sigmoid function can be computed in two ways as given in (21). Depending on the sign of $x$ it can be ensured that the argument of $\exp$ is always smaller than 00 to avoid numerical overflow. Therefore, we do not need to control the magnitude of $x$ externally with a max state and as a consequence also drop the normalizer state (see Appendix D.1). This yields the mLSTM with sigmoid input gate (henceforth referred to as mLSTMsig) in its recurrent formulation as
$\bm{C}_{t}$ $=\ \sigma(\tilde{\mathrm{f}}_{t})\ \bm{C}_{t-1}\ +\ \sigma(\tilde{\mathrm{i}}_{t})\ {\bm{k}_{t}\ \bm{v}_{t}^{\top}}$ (22)
@@
$\bm{h}_{t}$ $=\ \sigma\left(\tilde{\mathbf{o}}_{t}\right)\ \odot\ \mathrm{NORM}\left(\widetilde{\bm{h}}_{t}\right)$ (24)
where the query, key, and value vectors $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}$, and the gate preactivations $\tilde{\mathrm{i}}_{t},\tilde{\mathrm{f}}_{t},\tilde{\mathbf{o}}_{t}$ remain the same as for the mLSTM with exponential input gate (from now on referred to as mLSTMexp) in Section 2.1.
In Section 5.2, we confirm that our TFLA mLSTMsig forward kernel is over 30% faster than the mLSTMexp forward. We also show that mLSTMsig performs equally well compared to mLSTMexp in our language modeling experiments up to 1.4B parameters (see Section 5.1).
@@ 4.2 Normalization of mLSTM and Linear RNNs
We will see that the normalization layer $\bm{y}=\mathrm{NORM}(\bm{x})$, will play a crucial role in our analysis. The default norm layer in language modeling, the RMS norm (Zhang & Sennrich, 2019) with input vector input vector $\bm{x}\in\mathbb{R}^{d}$ and output vector $\bm{y}\in\mathbb{R}^{d}$ is defined as
$\bm{y}=\frac{\bm{x}}{\mathrm{RMS}(\bm{x})}\odot\bm{\gamma},\ \text{where}\ \mathrm{RMS}(\bm{x})=\sqrt{\frac{1}{d}\sum_{i=1}^{d}\ x_{i}^{2}+\epsilon},$ (25)
with with $\bm{\gamma}\in\mathbb{R}^{d}$ being a learnable scale parameter. The epsilon parameter $\epsilon\in\mathbb{R}$ is a small constant typically set to 1e-6 to avoid division by zero.
@@
where $\lVert\bm{x}\rVert_{\mathrm{max}}:=\max(|\bm{x}_{1}|,\dots,|\bm{x}_{d}|)$ and we average over the time dimension. For more details see App. D.2.
In Figure 4 we observe that the transfer behavior of mLSTMsig without normalizer is identical to mLSTMexp with normalizer and max state. Both exhibit a transition from suppressing ($G=0$) to passing ($G=1$) the signal at larger negative input gate preactivation values, which could partly explain the matching performance in our language modeling experiments.
Relation to other Gated Linear RNNs.
Interestingly, almost all other gated linear RNN variants also place a normalization layer after the RNN cell (Dao & Gu, 2024; Sun et al., 2023; Qin et al., 2024b; Yang et al., 2024b). Often this is justified with improved training stability, but a more thorough discussion is missing (Lieber et al., 2024). Qin et al. (2022) analyze the effect of the norm layer after a non-gated, kernel-based linear attention layer (Katharopoulos et al., 2020) and show that this effectively prevents unbounded gradients. We also confirm that the norm layer has a significant impact on training stability and the gradient norm during training. In Section 5.1 we show that initializing the input gate bias at larger negative values, as suggested by our transfer behavior analysis in Figure 4, prevents large gradient norm variance and spikes during training.
@@ 5.1 Language Modeling with mLSTM
We analyze the effect of the input gate bias initialization on training stability and performance of our mLSTM models in Appendix E.2. We observe in Figure 11 and 12, that initializing the input gate biases to -10 effectively mitigates large gradient norm spikes and variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 6), but also improves training stability, if initialized correctly.
@@
In Appendix E.2, we investigate the effect of the norm layer epsilon on language modeling performance for mLSTMexp. Our transfer behavior analysis in Figure 4 suggests, that there exists an interplay between norm layer epsilon and input gate bias initialization. We confirm this in our grid search in Figure 13 and find that the best performing configuration is the default epsilon ϵ=italic-$\epsilon=$1e-6 with input gate biases initialized to -10.
@@ 5.2 Kernel Benchmark
We compare the runtime of our mLSTM limit_chunk and TFLA xl_chunk kernels with kernel implementations of the state-of-the-art sequence modeling primitives FlashAttention (Dao, 2024; Shah et al., 2024), Mamba (Gu & Dao, 2024; Dao & Gu, 2024) and GLA (Yang et al., 2024b). In Appendix E.3 we compare with other kernels from the FlashLinearAttention library (Yang & Zhang, 2024). We run our benchmarks on NVIDIA H100 GPUs.
@@
In training, our TFLA kernels are faster than FlashAttention 3 for longer sequences and more than 2x faster than Mamba 2 kernels for all sequence lengths.
We perform additional runtime benchmarks for varying head dimensions and a more in-depth comparison to the FLA (Yang et al., 2024b) and LightningAttention2 (Qin et al., 2024a) kernels in Appendix E.3.
@@
The chunk size parameter $L$ balances the computation between the two levels of sequence parallelism (see Sec. 3.2). Smaller chunk sizes increase memory consumption, because more chunks are materialized in memory, but they reduce the quadratic compute FLOPs in the parallel part. Larger chunk sizes have the opposite effect. They decrease memory consumption, but increase quadratic compute FLOPs. In Figure 6, we measure this trade-off for our mLSTMsig TFLA xl_chunk kernels.
Figure 6: Memory vs. Runtime Trade-off of TFLA Forward-Backward Pass. We show the mLSTMsig for embedding dimension 4096 (8 heads with head dim 512), sequence length 8192 and batch size 8. By varying the chunk size parameter, our TFLA kernels can effectively balance memory vs. runtime.
@@ 5.3 Theoretical Runtime and Arithmetic Intensity
In Figure 6, we empirically observe that there exists an optimal chunk size (between 128 and 256) at which the runtime is minimized. In Appendix G, we compute the theoretical runtime optimal chunksize and the arithmetic intensity of TFLA depending on the chunk size by taking the FLOPs, memory operations and hardware accelerator specification into account.
We find that the theoretical runtime optimum exceeds the empirically observed one (see Figure 20 in Appendix G.1), suggesting that our current kernel implementations may not yet fully exploit the available performance potential.
We analyze the effect of the chunk size parameter $L$ on runtime, FLOPs, and arithmetic intensity in detail in Appendix F and G summarize our findings as follows:
  1. 1.
  2. 2.
  3. 3.
    [Equation (113), Figure 18, Figure 19]
  4. 4.
  5. 5.
    The chunk size determines whether the kernel is memory-bound or compute-bound on a given hardware. [Figure 22, Figure 21]
  6. 6.
    [Figure 23, Figure 20]
  7. 7.
    [Figure 24, Figure 25]
  8. 8.
@@ 6 Related Work
TFLA builds on concepts from FlashAttention (Dao, 2024) and Flash Linear Attention (Yang et al., 2024b), and is tailored for efficient mLSTM kernels (Beck et al., 2024). At the same time, it remains applicable to other linear RNNs such as RetNet, Mamba 2, and DeltaNet (Sun et al., 2023, 2024; Dao & Gu, 2024; Yang et al., 2024a). Additional discussion of these connections can be found in Appendix A.
7 Conclusion and Future Work
Although we enhance training stability through careful gate initialization informed by our empirical transfer behavior analysis, future work could explore instabilities arising from numerical errors in kernel implementations in greater depth. Finally, the programming techniques and hardware features used to optimize FlashAttention (Shah et al., 2024) could also be applied to our TFLA algorithm to approach peak performance on next-generation hardware, as suggested by our theoretical runtime analysis. This makes us believe that TFLA has the potential to become a foundational primitive for future long-context language models.
Acknowledgements
Impact Statement
The increasing adoption of neural networks in diverse fields, from healthcare to autonomous systems, has led to growing concerns about their environmental and economic costs. This paper presents novel techniques to enhance the energy and resource efficiency of neural networks. By reducing computational overhead and optimizing memory usage, our approach directly addresses the pressing need for sustainable AI practices.
While the proposed techniques improve the efficiency of neural networks, this benefit could be offset if practitioners leverage the increased efficiency to train larger models for longer durations, highlighting the need for conscientious application of these methods to truly achieve sustainable AI.
Appendix
[…]
       A.1. Relation to Flash Attention and Flash Linear Attention  A.1
       A.2. Other Related Work  A.2
       A.3. Application of TFLA to other Linear RNNs  A.3
   Appendix B - Extended mLSTM Formulations   B
       B.1. Fully Parallel Formulation  B.1
       B.2. Detailed Chunkwise-Parallel Formulation  B.2
       B.3. Chunkwise-Parallel Backward Pass  B.3
   Appendix C - Extended Tiled Flash Linear Attention   C
       C.1. Tiled Computation  C.1
       C.2. TFLA Forward Pass  C.2
       C.3. TFLA Backward Pass  C.3
   Appendix D - Extended mLSTM with Sigmoid Input Gate   D
       D.1. Stabilization of the Exponential Input Gate  D.1
       D.2. Empirical Transfer Behavior Analysis of the mLSTM  D.2
   Appendix E - Extended Experiments   E
       E.1. Numerical Validation of TFLA Kernels  E.1
       E.2. Extended Language Modeling Experiments with mLSTM  E.2
       E.3. Extended Kernel Benchmark  E.3
   Appendix F - FLOP and Memory Operation Counts for the mLSTM   F
       F.1. Exact FLOPs and Memory Operations Count for the mLSTM  F.1
       F.2. Simplified FLOP Count Summary for the mLSTM  F.2
       F.3. FLOP Comparison between mLSTMexp and mLSTMsig  F.3
       F.4. FLOP Comparison between different mLSTM Formulations  F.4
       F.5. FLOP-Optimal Chunk Size for mLSTM  F.5
   Appendix G - Theoretical Runtime Analysis for TFLA mLSTM Kernels   G
       G.1. Theoretical Runtime  G.1
       G.2. Arithmetic Intensity  G.2
       G.3. Roofline Analysis  G.3
       G.4. Runtime-Optimal Chunk Size  G.4
       G.5. FLOP and Theoretical Runtime Analysis Summary  G.5
Appendix A Related Work
A.1 Relation to Flash Attention and Flash Linear Attention
Tiled Flash Linear Attention (TFLA) combines the idea of tiling one sequence dimension the attention matrix for better work partitioning (Dao, 2024) with the idea of dividing the sequence into chunks (Yang et al., 2024b). These two ideas yield the two levels of sequence parallelism for TFLA.
@@ A.1 Relation to Flash Attention and Flash Linear Attention
TFLA is also IO-aware and parallelizes over one sequence dimension of the intra-chunk $\bm{Q}\bm{K}^{\top}$ matrix as the second level of sequence parallelism. New hardware features will also speed up future TFLA implementations.
@@
A.2 Other Related Work

### v3

@@ 4 Faster mLSTM with Sigmoid Input Gate
The mLSTM with exponential gating (i.e. exponential input gate) introduced by Beck et al. (2024) requires to compute and keep track of two additional states, the normalizer state $\bm{n}_{t}$ and max state $m_{t}$, as we show in Appendix D.1. Both will increase kernel runtime: The normalizer must be computed through summations, and tracking the max state throughout the tiled computation in TFLA (see Section 3 and C.2) prevents efficient fusing of loops within the kernel (see Appendix C.3).
Additionally, our analysis in Section 4.2 suggests to initialize the input gate biases at larger negative values (e.g. -10), such that the input gate pre-activations can grow slowly during training. We observe that most of these values stay below 0 during training (see Figure 15 in Appendix E). Therefore, we seek an alternative activation function which is similar to the exponential function in the negative range, but bounded in the positive range. This suggests to use the sigmoid function $\sigma(x)=\frac{1}{1+\exp(-x)}=\frac{\exp(x)}{\exp(x)+1}$, which converges to $\exp(x)$ for $x\rightarrow$-$\infty$ and $1$ for $x\rightarrow\infty$.
@@
The sigmoid function can be computed in two ways as given above.Depending on the sign of $x$ it can be ensured that the argument of $\exp$ is always smaller than 0 to avoid numerical overflow. Therefore, we do not need to control the magnitude of $x$ externally with a max state and as a consequence also drop the normalizer state (see Appendix D.1). This yields the mLSTM with sigmoid input gate (henceforth referred to as mLSTMsig) in its recurrent formulation as
$\bm{C}_{t}\$$=\ \sigma(\tilde{\mathrm{f}}_{t})\ \bm{C}_{t-1}\ +\ \sigma(\tilde{\mathrm{i}}_{t})\ {\bm{k}_{t}\ \bm{v}_{t}^{\top}}$ (13)
@@
$\bm{h}_{t}\$$=\ \sigma\left(\tilde{\mathbf{o}}_{t}\right)\ \odot\ \mathrm{NORM}\left(\widetilde{\bm{h}}_{t}\right)$ (15)
where the query, key, and value vectors $\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}$, and the gate preactivations $\tilde{\mathrm{i}}_{t},\tilde{\mathrm{f}}_{t},\tilde{\mathbf{o}}_{t}$ remain the same as for the mLSTM with exponential input gate (from now on referred to as mLSTMexp) in Section 2.1. We confirm that our TFLA mLSTMsig forward kernel is over 30% faster than the mLSTMexp forward (see Section 5.2), and show that mLSTMsig performs equally well compared to mLSTMexp in our language modeling experiments up to 1.4B parameters (see Section 5.1).
@@ 4.2 Normalization of mLSTM and Linear RNNs
Figure 4: Transfer behavior of the mLSTM before and after the RMS-norm layer ($\epsilon=$1e-6) for different input and forget gate values. The color shows the gain of the mLSTM defined in (16). After the norm layer mLSTMexp and mLSTMsig exhibit the same transfer behavior.
@@
We will see that the normalization layer $\bm{y}=\mathrm{NORM}(\bm{x})$, will play a crucial role in our analysis. The default norm layer in language modeling, the RMS norm (Zhang & Sennrich, 2019) with input vector input vector $\bm{x}\in\mathbb{R}^{d}$ and output vector $\bm{y}\in\mathbb{R}^{d}$ is defined as $\bm{y}=\frac{\bm{x}}{\mathrm{RMS}(\bm{x})}\odot\bm{\gamma}$, where $\mathrm{RMS}(\bm{x})=\sqrt{\frac{1}{d}\sum_{i=1}^{d}\ x_{i}^{2}+\epsilon}$, with with $\bm{\gamma}\in\mathbb{R}^{d}$ being a learnable scale parameter. The epsilon parameter $\epsilon\in\mathbb{R}$ is a small constant typically set to 1e-6 to avoid division by zero.
@@
where $\lVert\bm{x}\rVert_{\mathrm{max}}:=\max(|\bm{x}_{1}|,\dots,|\bm{x}_{d}|)$ and we average over the time dimension. For more details see App. D.2. In Figure 4 we observe that the transfer behavior of mLSTMsig without normalizer is identical to mLSTMexp with normalizer and max state. Both exhibit a transition from suppressing ($G=0$) to passing ($G=1$) the signal at larger negative input gate preactivation values, which could partly explain the matching performance in our language modeling experiments.
Normalization Layers in other Gated Linear RNNs.
Interestingly, almost all other gated linear RNN variants also place a normalization layer after the RNN cell (Sun et al., 2023; Dao & Gu, 2024; Qin et al., 2024b; Yang et al., 2024b). Often this is justified with improved training stability, but a more thorough discussion is missing (Lieber et al., 2024). Qin et al. (2022) analyze the effect of the norm layer after a non-gated, kernel-based linear attention layer (Katharopoulos et al., 2020) and show that this effectively prevents unbounded gradients. We also confirm that the norm layer has a significant impact on training stability and the gradient norm during training. In Section 5.1 we show that initializing the input gate bias at larger negative values, as suggested by our transfer behavior analysis in Figure 4, prevents large gradient norm variance and spikes during training. Relatedly, the general effect of layer normalization in the Transformer architecture has been investigated in several studies (Xiong et al., 2020; Zhu et al., 2025).
@@ 5.1 Language Modeling with mLSTM
We analyze the effect of the input gate bias initialization on training stability and performance of our mLSTM models in Appendix E.2. We observe in Figure 12 and 13, that initializing the input gate biases to -10 effectively mitigates large gradient norm spikes and variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 7), but also improves training stability, if initialized correctly.
@@
In Appendix E.2, we investigate the effect of the norm layer epsilon on language modeling performance for mLSTMexp. Our transfer behavior analysis in Figure 4 suggests, that there exists an interplay between norm layer epsilon and input gate bias initialization. We confirm this in our grid search in Figure 14 and find that the best performing configuration is the default epsilon $\epsilon=$1e-6 with input gate biases initialized to -10.
@@ 5.2 Kernel Benchmark
We compare the runtime of our mLSTM limit_chunk and TFLA xl_chunk kernels with kernel implementations of the state-of-the-art sequence modeling primitives FlashAttention (Dao, 2024; Shah et al., 2024), Mamba (Gu & Dao, 2024; Dao & Gu, 2024) and GLA Yang et al. (2024b). In Appendix E.3 we compare with other kernels from the Flash Linear Attention library (Yang & Zhang, 2024). We run our benchmarks on NVIDIA H100 GPUs.
@@
In training, our TFLA kernels are faster than FlashAttention 3 for longer sequences and more than 2x faster than Mamba 2 kernels for all sequence lengths. We perform additional runtime benchmarks for varying head dimensions and a more in-depth comparison to the FLA (Yang et al., 2024b) and LightningAttention2 (Qin et al., 2024a) kernels in Appendix E.3.
@@
Figure 6: Memory vs. Runtime Trade-off of TFLA Forward-Backward Pass. We show the mLSTMsig for embedding dimension 4096 (8 heads with dim 512), sequence length 8192 and batch size 8. By varying the chunk size parameter, TFLA kernels can effectively balance memory vs. runtime.
The chunk size parameter $L$ balances the computation between the two levels of sequence parallelism (see Sec. 3). Smaller chunk sizes increase memory consumption, because more chunks are materialized in memory, but they reduce the quadratic compute FLOPs in the parallel part. Larger chunk sizes have the opposite effect. They decrease memory consumption, but increase quadratic compute FLOPs. In Figure 6, we measure this trade-off for our mLSTMsig TFLA xl_chunk kernels.
@@ 5.3 Theoretical Runtime and Arithmetic Intensity
In Figure 6, we empirically observe that there exists an optimal chunk size (between 128 and 256) at which the runtime is minimized. In Appendix G, we compute the theoretical runtime optimal chunksize and the arithmetic intensity of TFLA depending on the chunk size by taking the FLOPs, memory operations and hardware accelerator specification into account. We find that the theoretical runtime optimum exceeds the empirically observed one (see Figure 21 in Appendix G.1), suggesting that our current kernel implementations may not yet fully exploit the available performance potential. We analyze the effect of the chunk size parameter $L$ on runtime, FLOPs, and arithmetic intensity in detail in Appendix F and G summarize our findings as follows:
(1) The chunk size $L$ mediates a trade-off between runtime and GPU memory usage [Figure 6]. (2) $L$ determines the total compute in FLOPs: $L=1$ matches the recurrent formulation, while $L=T$ matches the parallel one [Figure 19]. (3) There exists an optimal chunk size $L\in[1,T]$ that minimized the total FLOP count [Equation (103), Figure 19, Figure 20]. (4) Increasing $L$ raises the arithmetic intensity of TFLA kernels [Equation (109), Figure 22]. (5) The chunk size determines whether the kernel is memory-bound or compute-bound on a given hardware [Figure 23, Figure 22]. (6) FLOPs/s alone can be misleading; the optimal chunk size should be chosen based on total runtime [Figure 24, Figure 21]. (7) The runtime-optimal chunk size scales proportionally with the square root of the head dimension and the accelerator’s computational intensity [Figure 25, Figure 26]. (8) Newer hardware generations require larger chunk sizes to approach peak performance. [Figure 26, Figure 21].
@@ 6 Related Work
Tiled Flash Linear Attention (TFLA) integrates the concept of tiling along one sequence dimension of the attention matrix for improved work partitioning (Dao, 2024) with the strategy of dividing the sequence into chunks (Yang et al., 2024b), yielding two levels of sequence parallelism (see Figure 7).
@@
TFLA is IO-aware and parallelizes over one sequence dimension of the intra-chunk $\bm{Q}\bm{K}^{\top}$ matrix as the second level of sequence parallelism. New hardware features will also speed up future TFLA implementations.
@@
Figure 7: Tiled Flash Linear Attention (TFLA) combines Flash Linear Attention and Flash Attention 2. Flash Attention (middle) employs tiling across the head dimensions (green & light blue), while Flash Attention (right) uses tiling in the sequence dimension (orange & purple). Tiled Flash Linear Attention (left) combines both tiling strategies.
Application of TFLA to other Linear RNNs.
While TFLA is designed for efficient mLSTM kernels (Beck et al., 2024), its formulation also extends naturally to other linear RNNs such as RetNet, Mamba 2, and DeltaNet (Sun et al., 2023, 2024; Dao & Gu, 2024; Yang et al., 2024a). A more detailed discussion of these extensions is provided in Appendix A.
7 Conclusion and Future Work
Although we enhance training stability through careful gate initialization informed by our empirical transfer behavior analysis, future work could explore instabilities arising from numerical errors in kernel implementations in greater depth. Finally, the programming techniques and hardware features used to optimize Flash Attention (Shah et al., 2024) could also be applied to our TFLA algorithm to approach peak performance on next-generation hardware, as suggested by our theoretical runtime analysis. This establishes TFLA as a strong candidate for a foundational primitive in future long-context language models.
Acknowledgements
Appendix
  Appendix˜A - Extended Related Work   A
   A.1. Other Related Work . A.1
   A.2. Application of TFLA to other Linear RNNs . A.2
  Appendix˜B - Extended mLSTM Formulations   B
   B.1. Fully Parallel Formulation . B.1
   B.2. Detailed Chunkwise-Parallel Formulation . B.2
   B.3. Chunkwise-Parallel Backward Pass . B.3
  Appendix˜C - Extended Tiled Flash Linear Attention   C
   C.1. GPU Fundamentals . C.1
   C.2. Tiled Computation . C.2
   C.3. TFLA Forward Pass . C.3
   C.4. TFLA Backward Pass . C.4
  Appendix˜D - Extended mLSTM with Sigmoid Input Gate   D
   D.1. Stabilization of the Exponential Input Gate . D.1
   D.2. Empirical Transfer Behavior Analysis of the mLSTM . D.2
  Appendix˜E - Extended Experiments   E
   E.1. Numerical Validation of TFLA Kernels . E.1
   E.2. Extended Language Modeling Experiments with mLSTM . E.2
   E.3. Extended Kernel Benchmark . E.3
  Appendix˜F - FLOP and Memory Operation Counts for the mLSTM   F
   F.1. Exact FLOPs and Memory Operations Count for the mLSTM . F.1
   F.2. Simplified FLOP Count Summary for the mLSTM . F.2
   F.3. FLOP Comparison between mLSTMexp and mLSTMsig . F.3
   F.4. FLOP Comparison between different mLSTM Formulations . F.4
   F.5. FLOP-Optimal Chunk Size for mLSTM . F.5
  Appendix˜G - Theoretical Runtime Analysis for TFLA mLSTM Kernels   G
   G.1. Theoretical Runtime . G.1
   G.2. Arithmetic Intensity . G.2
   G.3. Roofline Analysis . G.3
   G.4. Runtime-Optimal Chunk Size . G.4
   G.5. FLOP and Theoretical Runtime Analysis Summary . G.5
Appendix A Extended Related Work
A.1 Other Related Work